An object starts from rest, accelerates at 1.2 m s-2
for 3.0 s and then moves at a
constant velocity for 6.0 s. Determine the final velocity and sketch the
velocity-time graph.
1
Expert's answer
2021-08-23T12:35:16-0400

We can find the final velocity from the kinematic equation:


"v=v_0+at=0+1.2\ \dfrac{m}{s^2}\cdot3.0\ s=3.6\ \dfrac{m}{s}."

Let's sketch the velocity-time graph: